# Fault Diagnosis and Prognosis Capabilities for Wind Turbine Hydraulic Pitch Systems: An Overview

### v1

@@ 3 Structural Analysis of the Benchmark Model
A mathematical model of a system can be translated into a set of constraints for it. The violation of some of them is an indication that a fault occurred, as the model is not able to describe the correct dynamics anymore. By combining the information of the constraints with the knowledge of the possible faults that may enter the plant, it is possible to understand which of the latter are detectable and, eventually, isolable, with the given plant configuration. It is worth remarking that these properties only depend on the system’s structure and not on the design of the diagnostic system. Thus, a structural analysis advises on possibilities for detectability and isolability of faults in a system based on the knowledge of its structure, i.e., its structural detectability and isolability properties.
@@
Structural analysis is a graph-based method to represent a system topology. The basis is a graph consisting of: a) function blocks, each of which represents constraints that describe behaviors of the function block; b) variables in the system; c) paths that describe links between constraints and variables. A system’s topology is described as a graph having constraints as the set of vertices a), variables in the system, b), and edges c) that connect the two types of vertices. The essence of structural analysis is to use tools from the arsenal of graph theory to obtain causal matchings in the bi-partitioned graph. A matching consists of a set of paths between unknown variables and constraints (dulmageCoveringsBipartiteGraphs1958, ). When an unknown variable $x$ (e.g. a state or a disturbance) is matched to a constraint $c$, the latter can be used for calculating the unknown variable. When all unknown variables are matched to constraints, any additional unmatched constraints can be used to describe redundant information in a system and constitute so called residuals. Residuals would be zero when there are no violations of normal behaviors and some would be non-zero if one or more behaviors were violated. Structural models were described at a tutorial level in (blankeStructuralDesignSystems2007, ) and references herein and software tools are available for swift analysis of structural properties. The SaTool (blankeSaToolSoftwareTool2006, ) focuses on violation of constraints and residual generation, including the computationally efficient method in (krysanderEfficientAlgorithmFinding2008a, ).
@@
Based on its graph-based representation, a system can be decomposed into three subsystems, i.e., three sub-graphs. The over-determined part contains unknown variables that are all matched as well as unmatched constraints. The just-determined subsystem contains equal number of constraints and unknown variables, which are matched to each other. Finally, the under-determined part contains unmatched unknown variables. A necessary condition for fault detection is that the over-constraint subsystem is nonempty and the under-constraint subsystem is the empty set. The Dulmage-Mendelson (DM) algorithm (dulmageStructureTheoryBipartite1959, ) is the original method for obtaining the structural decomposition of a system. Constraints are divided into just-determined and over-determined parts. If a fault appears in the just-determined part, no redundancy is present, and it is not detectable. The redundant equations in the over-determined part can be used for detecting faults that are affecting them. As a result, faults in the over-determined part are structurally detectable. Determining whether a fault is also isolable, requires a further analysis that includes looking into generation of residuals. An efficient way of determining the isolability properties was developed in (chenEfficientMethodDetermining2021, ).
@@ 3.1 Full Model Analysis
The analysis is carried out by referring to the equations stated in Section 2, with the support of the ’Fault Diagnosis Toolbox’ from (friskToolboxAnalysisDesign2017, ). The standard sensors present in a real system are utilized. The analysis hence delivers a benchmark result for the capabilities of diagnosis in such a plant. Different sensor configurations will be tested in Section 4 to obtain an equipment topology that maintains the diagnostic features of the original one, but with a reduced set of sensors. The Dulmage-Mendelson’s Decomposition for the standardized hydraulic pitch system, together with the partition in equivalent classes, is presented in Figure 3.
@@
The other faults that are group-wise isolable are related to the supply line, i.e., $f_{acc}$,$f_{Q_{ru}}$, and $f_{Q_{rel}}$. The oil leakages in the rotary unit and in the relief valve are acting on the same part of the system, i.e., the supply line. Thus, they are expected not to be distinguishable. The gas leakage in the accumulator is affecting Equation 17, through 16 and 21, on which both the other two faults are acting. The accumulator fault is a slow-varying fault. In practice, it can be detected and isolated by estimating the gas pre-charge pressure (linigerSignalBasedGasLeakage2017, ). As a consequence, that fault could be removed from the analysis and analyzed in this different framework. The other two faults can be isolated in practice by placing two pressure sensors, for example, before and after the rotary unit.
@@ 3.2 Distinct Operation Regions
1,2 9, 11 $x_{v}$ $x_{v}\lessgtr 0$
3,4 15 $p_{r}$, $p_{s}$ $p_{r}\lessgtr p_{s}-p_{cv,c}$
5,6 19 $p_{s}$ $p_{s}\lessgtr p_{cr,r}$
7,8 21 $p_{s}$ $p_{s}\lessgtr p_{gas,0}$
@@
The analysis of the separate operation regions follows the same procedure as the previous one, where the equations containing a Heaviside function are replaced with the one characterizing each specific case.
@@
The results are presented in the form of fault isolation matrices, as they represent the most straightforward tool for comparison, in Figure 6. The pattern of every plot coincides with every other and with the one in Figure 5. The outcome is that the analysis performed in Section 2 is meaningful for all the different conditions the plant is operating in. The full model can be utilized for deriving conclusions about sensor placement for improved diagnostic performance.
@@ 4 Sensor placement for topology improvement
In this section, different sensor combinations are investigated to determine whether it is possible to achieve the same standard model performance, by minimizing the number of sensors to be installed on the system. The tool that has been applied is still the Structural Analysis Toolbox (friskToolboxAnalysisDesign2017, ), by gradually adding sensors and evaluating the performance.
@@
For each of the three just-determined models, each possible set of sensors is added. The combination of $\{x_{c_{1}},x_{c_{2}},x_{c_{3}}\}$ with either $\{p_{p_{1}},p_{p_{2}},p_{p_{3}}\}$ or $\{p_{r_{1}},p_{r_{2}},p_{r_{3}}\}$ delivers isolability performances equivalent to the ones of the benchmark model. The combination of each of $\{p_{p_{1}},p_{p_{2}},p_{p_{3}}\}$ and $\{p_{r_{1}},p_{r_{2}},p_{r_{3}}\}$ with each other, or with any other set between $p_{s}$ and $\{x_{v_{1}},x_{v_{2}},x_{v_{3}}\}$, does not lead to any performance improvement. The same applies for the combination of $\{x_{c_{1}},x_{c_{2}},x_{c_{3}}\}$ with $p_{s}$ and/or $\{x_{v_{1}},x_{v_{2}},x_{v_{3}}\}$.
@@
$\{x_{c_{1}},x_{c_{2}},x_{c_{3}}\}$, $\{p_{p_{1}},p_{p_{2}},p_{p_{3}}\}$
Cylinder pistons positions, all piston-side pressures
$\{x_{c_{1}},x_{c_{2}},x_{c_{3}}\}$, $\{p_{p_{1}},p_{p_{2}}\}$, $\{p_{r_{3}}\}$
Cylinder pistons positions, piston-side and rod-side pressures
$\{x_{c_{1}},x_{c_{2}},x_{c_{3}}\}$, $\{p_{p_{1}}\}$, $\{p_{r_{2}},p_{r_{3}}\}$
Cylinder pistons positions, piston-side and rod-side pressures
$\{x_{c_{1}},x_{c_{2}},x_{c_{3}}\}$, $\{p_{p_{1}},p_{p_{3}}\}$, $\{p_{r_{2}}\}$
Cylinder pistons positions, piston-side and rod-side pressures
$\{x_{c_{1}},x_{c_{2}},x_{c_{3}}\}$, $\{p_{p_{2}}\}$, $\{p_{r_{1}},p_{r_{3}}\}$
Cylinder pistons positions, piston-side and rod-side pressures
$\{x_{c_{1}},x_{c_{2}},x_{c_{3}}\}$, $\{p_{p_{2}},p_{p_{3}}\}$, $\{p_{r_{1}}\}$
Cylinder pistons positions, piston-side and rod-side pressures
$\{x_{c_{1}},x_{c_{2}},x_{c_{3}}\}$, $\{p_{p_{3}}\}$, $\{p_{r_{1}},p_{r_{2}}\}$
Cylinder pistons positions, piston-side and rod-side pressures
$\{x_{c_{1}},x_{c_{2}},x_{c_{3}}\}$, $\{p_{r_{1}},p_{r_{2}},p_{r_{3}}\}$
Cylinder pistons positions, all rod-side pressures

### v2

@@ 3 Structural Analysis of the Benchmark Model
This section provides an introduction to the key structural properties of dynamical systems that are used for fault detection and presents a systematic assessment of the system’s diagnostic capabilities. The structural analysis is first performed at a generic operating point and later separately applied to all the individual operating regions.
3.1 Analysis based on system’s structure
The modelling needed for fault-tolerant design pertains to the topology of the system and is based on the individual relations between the variables. This is done through formulation of constraints. When modelling to obtain overall functional relations, it is desirable not to predefine what is input and what is output of individual function blocks, but rather state that there are relations between these variables, constrained by given relations. Such constraints were described in the modelling Sections 2.1 to 2.3. Modelling for fault-tolerance is therefore conveniently done using the principles of behavioural modelling, where constraints $c$ describe how variables are related. Let variables be $x,z$ where $x$ are unknown, $z$ are known (input $u$ or output $y$), and disturbances are $d$. Let $g_{s}$ and $g_{d}$ denote functions; then constrains can be static $(c_{s})$ or dynamic $(c_{d})$ :
$\begin{array}[]{c c c c}c_{s}:&0&=&g_{s}(x,d)-z\\ c_{d}:&0&=&g_{d}(x,d)-\dot{x},\\ \end{array}$
where $g_{s}$ and $g_{d}$ can be linear or nonlinear. Derivatives of variables can be explicit or implicit in the constraints.
@@ 3.1 Analysis based on system’s structure
A violation of one or more constraints is an indication that some fault has occurred, as the model is not able to describe the correct dynamics anymore. By combining the information of the constraints with the knowledge of the possible faults that may enter the plant, it is possible to understand which of the latter are detectable and, eventually, isolable, for a given plant configuration (e.g. sensor placement list). It is worth noting that these properties only depend on the system’s structure and not on the design of the diagnostic system. Thus, a structural analysis advises on possibilities for detectability and isolability of faults in a system based on the knowledge of its structure. This provides us with its generic structural detectability and isolability properties.
@@
Structural analysis is a method where a graph represents system topology. The basis is a graph consisting of: 1) function blocks, each of which represents constraints that describe behaviors of the function block; 2) variables in the system; 3) paths that describe links between constraints and variables. A system’s topology is described as a graph having constraints as the set of vertices a), variables in the system, b), and edges c) that connect the two types of vertices. The essence of structural analysis is to use tools from the arsenal of graph theory to obtain causal matchings in the bi-partitioned graph. A matching consists of a set of paths between unknown variables and constraints [18]. When an unknown variable $x$ (e.g. a state or a disturbance) is matched to a constraint $c$, the latter can be used to calculate the unknown variable. When all unknown variables are matched to constraints, any additional unmatched constraint can be used to describe redundant information to compute so called residuals. Residuals would be zero when there are no violations of normal behaviors and some will be non-zero if one or more behaviors are violated. Structural models were described at a tutorial level in [7] and references herein and software tools are available for swift analysis of structural properties. The SaTool [6] focuses on violation of constraints and residual generation, including the computationally efficient method in [32].
@@
Based on its graph-based representation, a system can be decomposed into three subsystems, i.e., three sub-graphs. The over-determined part contains unknown variables that are all matched as well as unmatched constraints. The just-determined subsystem contains equal number of constraints and unknown variables, which are matched to each other. Finally, the under-determined part contains unmatched unknown variables. A necessary condition for fault detection is that the over-constraint subsystem is nonempty and the under-constraint subsystem is the empty set. The Dulmage-Mendelson (DM) algorithm [17] is the original method for obtaining the structural decomposition of a system. Constraints are divided into just-determined and over-determined parts. If a fault appears in the just-determined part, no redundancy is present, and it is not detectable. The redundant equations in the over-determined part can be used for detecting faults that are affecting them. As a result, faults in the over-determined part are structurally detectable. Determining whether a fault is also isolable, requires a further analysis that includes looking into generation of residuals. An efficient way of determining the isolability properties was developed in [9].
Analytical redundancy relations are found by using the minimal structural over-determined subsets (MSO) and by backtracking through a selected matching, $\mathcal{M}$. Starting with $c_{n+1}$ being an unmatched constraint, whereas $c_{n}$ to $c_{1}$ are matched, $c_{1},...,c_{n}\in\mathcal{M}_{N}$, then, with $z_{i}\in\mathcal{Z}$ where $\mathcal{Z}=\{x_{i},d_{j},f_{k}\}$ is the set of all unknown variables $x_{i}$, disturbances $d_{j}$, explicitly modelled faults, $f_{k}$, known input $u_{n}$, and measurements $y_{m}$,
$c_{N+1}(z_{1},..z_{N})=c_{N+1}(c_{1}(\{x_{i},d_{j},f_{k},u_{n},y_{m}\}\in\mathcal{Z}_{1},$ (23)
$...,c_{N}(\{x_{i},d_{j},f_{k},u_{n},y_{m}\}\in\mathcal{Z}_{N}),$
where the set $\mathcal{Z}_{1}\subset\mathcal{Z}$ is the parts of the sub-graph that are employed to match $c_{1}$, etc.
A residual $r$ is obtained from $c_{N+1}$ by iteration through the matching $\mathcal{M}_{1}$ that calculates the variables $x,d,f$ such that the residual becomes a function solely of the known variables $u,y\in\mathcal{M}_{1}$, i.e., a residual is,
$r(t)=g(\{c(\cdot),u(t),y(t)\}\in\mathcal{M}_{1}),$ (24)
which is evaluated in real time.
This construction of residuals from structural analysis guarantees that violation of any constraints are structurally detectable. It does not express the sensitivity with which a particular residual can reflect a violation of constraints, i.e. fault(s) in the system since sensitivity is not a structural property, but a property in the analytic domain.
Design of diagnosis and monitoring algorithms therefore has two steps. One is analysis of structure that shows which possibilities of change detection exist with a given system topology. The second step is to materialize this in the analytic domain where sensitivity to changes (faults) and robustness to disturbances and parameter uncertainty are addressed.
@@ 3.2 Full Model Analysis
The analysis is carried out by referring to the equations stated in Section 2, with the support of the ’Fault Diagnosis Toolbox’ from [23]. The standard sensors present in a real system are utilized. The analysis hence delivers a benchmark result for the capabilities of diagnosis in such a plant. Different sensor configurations will be tested in Section 4 to obtain an equipment topology that maintains the diagnostic features of the original one, but with a reduced set of sensors. The Dulmage-Mendelson’s Decomposition for the standardized hydraulic pitch system, together with the partition in equivalent classes, is presented in Figure 3.
@@
The other faults that are group-wise isolable are related to the supply line, i.e., $f_{acc}$,$f_{Q_{ru}}$, and $f_{Q_{rel}}$. The oil leakages in the rotary unit and in the relief valve are acting on the same part of the system, i.e., the supply line. Thus, they are expected not to be distinguishable. The gas leakage in the accumulator is affecting Equation 18, through 17 and 22, on which both the other two faults are acting. The accumulator fault is a slow-varying fault. In practice, it can be detected and isolated by estimating the gas pre-charge pressure [38]. As a consequence, that fault could be removed from the analysis and analyzed in this different framework. The other two faults can be isolated in practice by placing two pressure sensors, for example, before and after the rotary unit.
@@
The simplification in the effective bulk modulus affects the structure of the system, by lowering the number of unknown parameters in the model. The choice improves the detectability/isolability properties of the system, thanks to the higher number of redundant relation that are consequent. As anticipated in the previous section, the price to pay is the impossibility of distinguishing the exact location where the oil is degrading.
Since no substantial unmodelled dynamics have been reported in the well-established literature for the hydraulic actuator system, the dominant source of uncertainty in the model is associated with variations in the parameters that appear in its equations. Although parametric uncertainty may affect the performance of the residuals that can be derived based on the result of the structural analysis, the analysis itself, and hence, the diagnostic capabilities of the system, are invariant to such model mismatch. This is due to the fact that such analysis relies on the structural properties of the constraints. The robustness of the presented algorithms with respect to parametric uncertainty is discussed later in Section 5.
@@ 3.3 Distinct Operation Regions
1,2 10, 12 $x_{v}$ $x_{v}\lessgtr 0$
3,4 16 $p_{r}$, $p_{s}$ $p_{r}\lessgtr p_{s}-p_{cv,c}$
5,6 20 $p_{s}$ $p_{s}\lessgtr p_{cr,r}$
7,8 22 $p_{s}$ $p_{s}\lessgtr p_{gas,0}$
@@
The analysis of the separate operation regions follows the same procedure as the previous one, where the equations containing a Heaviside function are replaced with the one characterizing each specific case. The results are presented in the form of fault isolation matrices, as they represent the most straightforward tool for comparison, in Figure 6. The pattern of every plot coincides with every other and with the one in Figure 5. The outcome is that the analysis performed in Section 2 is meaningful for all the different conditions the plant is operating in. The full model can be utilized for deriving conclusions about sensor placement for improved diagnostic performance.
@@ 4 Sensor placement for topology improvement
In this section, different sensor combinations are investigated to determine whether it is possible to achieve the same standard model performance, by minimizing the number of sensors to be installed on the system. The tool that has been applied is still the Structural Analysis Toolbox [23], by gradually adding sensors and evaluating the performance.
@@
For each of the three just-determined models, each possible set of sensors is added. The combination of $\{x_{c_{1}},x_{c_{2}},x_{c_{3}}\}$ with either $\{p_{p_{1}},p_{p_{2}},p_{p_{3}}\}$ or $\{p_{r_{1}},p_{r_{2}},p_{r_{3}}\}$ delivers isolability performances equivalent to the ones of the benchmark model.
The combination of each of $\{p_{p_{1}},p_{p_{2}},p_{p_{3}}\}$ and $\{p_{r_{1}},p_{r_{2}},p_{r_{3}}\}$ with each other, or with any other set between $p_{s}$ and $\{x_{v_{1}},x_{v_{2}},x_{v_{3}}\}$, does not lead to any performance improvement. The same applies for the combination of $\{x_{c_{1}},x_{c_{2}},x_{c_{3}}\}$ with $p_{s}$ and/or $\{x_{v_{1}},x_{v_{2}},x_{v_{3}}\}$.
@@
$\{x_{c_{1}},x_{c_{2}},x_{c_{3}}\}$, $\{p_{p_{1}},p_{p_{2}},p_{p_{3}}\}$ Cylinder pistons positions, all piston-side pressures
$\{x_{c_{1}},x_{c_{2}},x_{c_{3}}\}$, $\{p_{p_{1}},p_{p_{2}}\}$, $\{p_{r_{3}}\}$ Cylinder pistons positions, piston-side and rod-side pressures
$\{x_{c_{1}},x_{c_{2}},x_{c_{3}}\}$, $\{p_{p_{1}}\}$, $\{p_{r_{2}},p_{r_{3}}\}$ Cylinder pistons positions, piston-side and rod-side pressures
$\{x_{c_{1}},x_{c_{2}},x_{c_{3}}\}$, $\{p_{p_{1}},p_{p_{3}}\}$, $\{p_{r_{2}}\}$ Cylinder pistons positions, piston-side and rod-side pressures
$\{x_{c_{1}},x_{c_{2}},x_{c_{3}}\}$, $\{p_{p_{2}}\}$, $\{p_{r_{1}},p_{r_{3}}\}$ Cylinder pistons positions, piston-side and rod-side pressures
$\{x_{c_{1}},x_{c_{2}},x_{c_{3}}\}$, $\{p_{p_{2}},p_{p_{3}}\}$, $\{p_{r_{1}}\}$ Cylinder pistons positions, piston-side and rod-side pressures
$\{x_{c_{1}},x_{c_{2}},x_{c_{3}}\}$, $\{p_{p_{3}}\}$, $\{p_{r_{1}},p_{r_{2}}\}$ Cylinder pistons positions, piston-side and rod-side pressures
$\{x_{c_{1}},x_{c_{2}},x_{c_{3}}\}$, $\{p_{r_{1}},p_{r_{2}},p_{r_{3}}\}$ Cylinder pistons positions, all rod-side pressures
5 Robustness to model uncertainty
The main source of model uncertainty is related to the model parameters. Parameter values and uncertainty ranges are shown in Table 4. Additionally, the function $K(x_{v})$ in Equations 10 and 12 has an uncertainty of $5\%$.
[…]
$M_{eq}$ Cylinder’s piston equivalent mass 104 $Kg$ 1
$A_{p}$ Piston side cylinder’s cross section 0.015 $m^{2}$ 1
$A_{r}$ Rod side cylinder’s cross section 0.009 $m^{2}$ 1
$B_{v}$ Viscous friction coefficient 6500 $Nms^{-1}$ 80
$F_{c}$ Coulomb friction coefficient 2600 $N$ 80
$\gamma$ Sign function approximation $2\times 10^{-4}$- 0
$\beta_{e,0}$ Oil’s effective bulk modulus $7\times 10^{8}$ $Pa$ 40
$V_{0,p}$ Piston side dead volume $3.7\times 10^{-4}$ $m^{3}$ 1
$V_{0,r}$ Rod side dead volume $1.5\times 10^{-4}$ $m^{3}$ 1
$x_{c,max}$ Cylinder’s stroke 1.35 $m$ 1
$C_{le,p}$ External leakage coefficient (piston) 0 $m^{3}sKg^{-1}$ 0
$C_{le,r}$ External leakage coefficient (rod) 0 $m^{3}sKg^{-1}$ 0
$C_{li}$ Internal leakage coefficient 0 $m^{3}sKg^{-1}$ 0
$\phi_{v}$ Cross sections ratio 0.5867 -- 5
$\xi$ Valve’s Damping ratio 1 -- 1
$\omega_{0}$ Valve’s Natural frequency 440 $rads^{-1}$ 1
$K_{cv}$ Check valve flow coefficient $1.5\times 10^{-8}$ $m^{3}sKg^{-1}$ 10
$p_{cv,c}$ Check valve crack pressure $0.5\times 10^{5}$ $Pa$ 50
$k$ Accumulator’s gas stiffness 1.4 $m^{3}sKg^{-1}$ 0
$K_{rel}$ Relief valve flow coefficient $1.5\times 10^{-8}$ $m^{3}sKg^{-1}$ 10
$p_{cr,r}$ Relief pressure $2.1\times 10^{7}$ $Pa$ 0.5
$V_{acc}$ Accumulator volume 0.1 $m^{3}$ 1
$V_{hose}$ Hose volume $1\times 10^{-3}$ $m^{3}$ 1
$p_{gas,0}$ Accumulator nominal pressure $1.89\times 10^{7}$ $Pa$ 0.5
Table 4: Table of parameters with their values and uncertainty ranges (in percentage).
5.1 Model uncertainty
Denote the modelled behaviour by $\hat{g}$, and the actual by $\breve{g}$, then modelled and actual residuals become parameterised in $g$ as,
$\hat{r}(t)$ $=\hat{g}(\{\hat{c}(\cdot),u(t),y(t)\}\in{\mathcal{M}}_{1})$ (25)
$\breve{r}(t)$ $=\breve{g}(\{{c}(\cdot),d(t),f(t),u(t),y(t)\}\in{\mathcal{M}}_{1})$ (26)
where $d$ represent disturbances, $u$ is input, $y$ the measurements and $f$ the signals representing faults. It should be noted that in the absence of faults and disturbances the actual and model residual signals are equal, i.e. $\breve{r}(t)=\hat{r}(t)$ when $d(t)=0,\;f(t)=0$.
Perturbation analysis then provides the sensitivity of the residual to uncertainty of parameters, i.e. to a variation of the parameter vector $\Delta p=\breve{p}-\hat{p}$. Indeed, using Taylor expansion around the values used in the model for the parameters, disturbances and the considered fault, the actual residual $\breve{r}$ is approximated as:
$\breve{r}(t)$ $\approx\underbrace{\frac{\partial\breve{r}}{\partial y}\Delta y+\frac{\partial\breve{r}}{\partial u}\Delta u}_{\hat{r}(t)}+\underbrace{\frac{\partial\breve{r}}{\partial p}\Delta p+\frac{\partial\breve{r}}{\partial d}d+\frac{\partial\breve{r}}{\partial f}f}_{\tilde{r}(t)\triangleq\breve{r}(t)-\hat{r}(t)}$ (27)
and the consequences of model uncertainty may include:
  • 1.
    the residual becomes sensitive to disturbances, whereas, by design of $r$, the signals $d$, $u$ and $y$ do not affect the residual in the nominal case.
  • 2.
    some faults $f$ may be masked by $\Delta p$ and $d$.
The latter implies that when evaluating $\hat{r}(t)$ by comparing single samples with a threshold $T$ in the noise free case, safe conclusions on the detection of a fault $f$ can be drawn if the following condition holds
$\|f\|_{\infty}>\frac{\left\|\displaystyle\frac{\partial\breve{r}}{\partial p}\right\|_{\infty}\|\Delta p\|_{2}+\left\|\displaystyle\frac{\partial\breve{r}}{\partial d}\right\|_{\infty}\|d\|_{\infty}}{\left\|\displaystyle\frac{\partial\breve{r}}{\partial f}\right\|_{\infty}}$ (28)
where the $\mathcal{L}_{\infty}$ norm of a time-varying vector (equivalently matrix) $y(t)$ is defined as $\|y\|_{\infty}\triangleq\sup\limits_{t}\|y(t)\|$ using any vector (equivalently induced matrix) norm.
5.2 Example
As an example, a residual associated with the results from the structural analysis in Section 3 is given in analytical form by
$\breve{r}=$ $-\frac{\dot{p}_{r}}{B_{e}}[V_{or}-A_{r}(x_{c}-x_{c,max})]+A_{r}\dot{x}_{c}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}-f_{Qle,r}}$
$+\phi_{v}K_{v}(x_{v})\sqrt{|p_{r}-p_{s}|}\textit{sign}(p_{r}-p_{s})\textit{H}(-x_{v})$
$+K_{cv}(p_{s}-p_{r}-p_{cv_{c}})\textit{H}(p_{s}-p_{r}-p_{cv_{c}})$ (29)
Consider an external leakage fault $f=f_{Qle,r}$. The bulk modulus $B_{e}$ and the check valve flow coefficient $K_{cv}$ are uncertain parameters with variations $|\Delta B_{e}|<\delta_{1}$ and $|\Delta K_{cv}|<\delta_{2}$. The bounds $\delta_{1},\delta_{2}>0$ are known from Table 4. Following the notation introduced in (28) with $p\triangleq\begin{bmatrix}B_{e}&K_{cv}\end{bmatrix}^{T}$ and $\Delta p\triangleq\begin{bmatrix}\Delta B_{e}&\Delta K_{cv}\end{bmatrix}^{T}$ one gets
$\frac{\partial\breve{r}}{\partial p}$ $=\begin{bmatrix}\displaystyle\frac{\dot{p}_{r}}{B_{e}^{2}}[V_{or}-A_{r}(x_{c}-x_{c,max})]&(p_{s}-p_{r}-p_{cv_{c}})\textit{H}(p_{s}-p_{r}-p_{cv_{c}})\end{bmatrix}$
$\frac{\partial\breve{r}}{\partial d}$ $=0$ (30)
$\frac{\partial\breve{r}}{\partial f_{Qle,r}}$ $=-1\;.$
Let $\phi>0$ be a bound on $\left\|\frac{\partial\breve{r}}{\partial p}\right\|_{\infty}$, i.e.
$\sup\limits_{t}\sqrt{\displaystyle\frac{\dot{p}^{2}_{r}}{B_{e}^{4}}[V_{or}-A_{r}(x_{c}-x_{c,max})]^{2}+(p_{s}-p_{r}-p_{cv_{c}})^{2}}\leq\phi.$ (31)
Then application of condition (28) leads to the requirement
$\|f_{Qle,r}\|_{\infty}>\phi\sqrt{\delta_{1}^{2}+\delta_{2}^{2}}\;.$ (32)
The presence of a fault is detected if $|\hat{r}(t)|>\phi$, which ensures that the detection of the fault is robust to parameter variations. Estimating $\phi$ is subject to the values obtained by the states and outputs of the system, i.e., to excitation conditions and operating mode. This suggests that such bounds can often be conservative even if $\delta_{1},\delta_{2}$ are not.